OpenStudy (anonymous):

Grouping Method! 5x^2-25x-60

OpenStudy (anonymous):

My teacher only wants the grouping method... thats the problem.

OpenStudy (anonymous):

Yes, but it doesnt fit for this equation... Ill post it.

OpenStudy (anonymous):

After Grouping : \[[x -( 5-\sqrt{73}/2)][x -(5+\sqrt{73})/2] = 0 \] But this method is inconvenient for this question!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

You cannot use grouping method for this. You can use it only if it +60 not -60.

OpenStudy (anonymous):

Use Feynman Algorithm : Write down the problem,Try real hard,Write down the solution